# Hawaii: Hierarchical Visual Knowledge Transfer for Efficient Vision-Language Models

### v1

@@ A.1 Benchmark Datasets
A.2 Ablation Study
[…]
[width=0.32]pics/routing/expert_choice_AI2D_v1.0.pdf
[…]
[width=0.32]pics/routing/expert_choice_GQA_v1.0.pdf
[…]
[width=0.32]pics/routing/expert_choice_MMBench_v1.0.pdf
[…]
[width=0.32]pics/routing/expert_choice_MME_v1.0.pdf
[…]
[width=0.32]pics/routing/expert_choice_MMMU_v1.0.pdf
[…]
[width=0.32]pics/routing/expert_choice_POPE_v1.0.pdf
[…]
[width=0.32]pics/routing/expert_choice_SEEDBench_v1.0.pdf
[…]
[width=0.32]pics/routing/expert_choice_TextVQA_v1.0.pdf
[…]
[width=0.32]pics/routing/expert_choice_VizWiz_v1.0.pdf
@@ A.2 Ablation Study
Routing between specific teachers’ knowledge. To further understand how Hawaii switches between different teachers’ knowledge, we visualize the routing results in Figure 6. It is obvious that Hawaii selects different expert’s knowledge across different benchmark datasets and different layers. A notable observation is for most of the cases, Hawaii does not choose CLIP for understanding visual contents. We observe that for MME, VizWiz, and SEEDBench, the model has similar selection preference, while for MMMU, model mainly choose CLIP and ConvNext.

### v2

@@ A.1 Benchmark Datasets
A.2 Comparison with MLLMs with Multiple Vision Encoders
To better understand how Hawaii compares with the existing MLLMs with multiple vision encoders [13, 14, 47, 68], we present the comparison in Table˜6. Results show that Hawaii achieves competitive or significant improvements on most benchmarks, demonstrating the effectiveness of Hawaii. However, we also notice performance degradation on some benchmarks, such as POPE, GQA, and SeedBench.
VizWiz GQA SQA POPE MME AI2D MMMU SeedBench
Eagle-X5 [47] 54.4 64.9 69.8 88.8 1528 - 36.3 73.9
MoME [14] (CLIP + DINO + Pix2Struct) - 59.7 - - - - - -
MouSi [68] (LayoutLMv3+DINOv2+CLIP) - 63.6 69.0 86.5 - - - 67.5
Brave [13] 54.2 52.7 - 87.6 - - - -
LLaVA-1.5 (CLIP) 50.0 62.0 66.8 85.9 1510.7 55.5 34.7 66.1
MoVE-KD 52.3 63.2 69.4 86.9 1524.5 - - -
HAWAII 53.9 62.8 70.5 87.3 1540.2 56.2 36.6 67.5
HAWAII† (HAWAII + Pix2Struct) 54.2 63.6 69.5 86.8 1533.6 56.2 36.0 67.2
HAWAII‡ (HAWAII + SAM) 54.3 63.3 70.8 87.5 1528.6 55.8 36.9 67.9
$\Delta$ $\downarrow$ 0.1 $\downarrow$ 1.3 $\uparrow$ 1.0 $\downarrow$ 1.3 $\uparrow$ 12.2 - $\uparrow$ 0.6 $\downarrow$ 6.0
Table 6: Comparison with MLLMs with multiple vision encoders.
A.3 Ablation Study
@@
Routing between specific teachers’ knowledge. To further understand how Hawaii switches between different teachers’ knowledge, we visualize the routing results in Figure˜6. It is obvious that Hawaii selects different expert’s knowledge across different benchmark datasets and different layers. A notable observation is for most of the cases, Hawaii does not choose CLIP for understanding visual contents. We observe that for MME, VizWiz, and SEEDBench, the model has similar selection preference, while for MMMU, model mainly choose CLIP and ConvNext.